OpenStudy (anonymous):

Solve using the substitution method 3x+2y=38 y=x+4

OpenStudy (anonymous):

y = -3/2x + 19 That's the slope of that line.

OpenStudy (anonymous):

3x + 2y = 38 3x + 2(x + 4)= 38 3x + 2x + 8 = 38 5x + 8 = 38 Subtract 8 from both sides 5x = 38-8 5x = 30 x = 30/5 or 6 Plug the x value into equation 1 to solve for y y = 6 + 4 y = 10 x=6 and y=10 or (6,10) this is what I got

OpenStudy (anonymous):

Anwser your own question. Can't help you here buhh bye.

OpenStudy (anonymous):

I only know how to do slope.

OpenStudy (anonymous):

If you only know how to do slope then why would you even attempt to answer it when it says solve using the substitution method

OpenStudy (anonymous):

Don't get a attitude k? And i thought you wanted slope.

OpenStudy (anonymous):

I don't have an attitude and if I wanted the slope I would have asked for it. I think you are rude for telling me to answer my own question when I was only looking for help to make sure I got it correct. That is what this is supposed to be is a helping site.